OpenStudy (anonymous):

I know to find the perimeter you have to add all side lengths but how would you find the perimeter if this triangle?

OpenStudy (anonymous):

.

OpenStudy (usukidoll):

perimeter of the triangle is P = a+b+c where: a,b and c are the lengths of each side of the triangle

OpenStudy (nuttyliaczar):

Add the sides like you normally would. 5w + (-2w +15) + (w - 7)

OpenStudy (nuttyliaczar):

And combine like terms

OpenStudy (usukidoll):

so let a = (w-7) b = (-2w+15) c = (5w) P = (w-7)+(-2w+15)+5w and combine like terms... wow mine is backwards. I wonder if order matters ?

OpenStudy (triciaal):

no order does not matter

OpenStudy (madhu.mukherjee.946):

[5w+w-7-2w+15]in =[4w+8]in